# Will anyone review this paper? Screening, sorting, and the feedback cycles that imperil peer review

### v2

@@ III.3.2 Several competing journals
A formal model that includes desk rejection with several competing journals lies outside the present scope, since it would require modeling how authors behave when they face journals with different desk-rejection cutoffs. However, intuition suggests that when competing journals rely on a common pool of reviewers, a “tragedy of the reviewer commons” [34] makes desk rejection less useful for managing the review load. The logic goes as follows. When a journal operates with its own private community of reviewers, it fully internalizes the cost of sending a manuscript our for review. At equilibrium, such a journal sets its desk-rejection policy to optimally trade off the cost of depleting its available review labor against the benefit of using peer review to learn more about a manuscript’s quality. But when several competing journals rely on the same common pool of reviewers, then a journal seeking a review does not fully internalize the cost of requesting that review, because the cost of depleting the available review labor is borne by all journals. Yet the requesting journal still captures the full benefit of obtaining a review and learning more about the manuscript under consideration. Thus, because journals that rely on a shared reviewer pool do not fully internalize the cost of requesting reviews, they will set a desk-rejection policy that desk-rejects manuscripts less often than would be socially optimal.
@@ IV Discussion
Like all models, our model makes simplifications that provide scope for future work. Perhaps the most substantively, we treat manuscript quality as exogenous. Of course, authors are not endowed with fully formed manuscripts, but instead they decide how much effort to invest to developing manuscripts based on their anticipation of the journal’s scrutiny. For example, the rise of revise-and-resubmit as an author’s likely best outcome may have the perverse effect of encouraging authors to submit manuscripts that are less polished, both because revision seems inevitable and because journals can no longer expect polished submissions if unpolished submissions become the norm [40]. A model of journal and author behavior that endogenizes manuscript quality would provide more satisfying insight on this count.
@@
Adverse selection for weaker manuscripts also explains why top journals are unlikely to simply tolerate a decline in review quality. Top authors benefit from accurate reviewing and thus prefer to submit their work to journals that offer it. Journals who seek to attract top authors must try to keep their reviewing accurate, forcing them to lean heavily on the review pool as the review burden increases. Again, this seems to describe the current publishing environment, where top journals typically seek more reviews than less prestigious ones.
@@
If increasing submission costs and tolerating less accurate reviews are off the table, journals might instead try to increase the supply of high-quality review labor. But journals that seek to recruit more high-quality reviewing are handcuffed by the tradition of peer review as an volunteer activity. In a typical labor market, employers can attract more labor by increasing wages. Journals operate without this basic market mechanism when reviewers are unpaid. If publishers were able to recruit more high-quality review labor by paying appropriate wages, they could maintain constant review quality in the face of an increase in submissions, breaking the feedback loop between screening and sorting.
@@
To be sure, paying reviewers brings risks. Non-profit journals would need to pass the cost of reviewer pay on to authors or subscribers, although perhaps society journals could maintain a base of volunteer reviewers by cultivating a sense of shared responsibility for the journal. Paying reviewers also commodifies review labor, likely irreversibly [28]. This commodification could perversely decrease the availability of high-quality review labor if reviewers operating under the honor system perceive their effort to be worth more than the wages that the journal offers [28, 29, 27]. Nevertheless, recent experiments with modest payments to reviewers have succeeded in accelerating the pace of peer review, at least in the short run [15, 16, 29]. While substantial trial and error should be expected in establishing a wage scale that adequately compensates reviewers, once such a wage scale is established, it could give publishers a lever to recruit reviewers that they currently lack.
@@
Are there other options that might salvage the institution of (voluntary) peer review? We highlight two options here, both of which echo the suggestions of others. First, journals can reduce re-review of revised manuscripts [21, 30]. Re-review is currently commonplace: accepted papers frequently go through at least two [35] or more [30] rounds of review prior to acceptance. But routinely sending revised manuscripts back out for review only burdens the reviewer community further while encouraging both authors and reviewers to modulate their effort in anticipation of the revision and re-review process to follow [40]. Journals can reduce re-review by empowering, or even encouraging, academic editors to be more proactive in their decision-making, especially when evaluating revised manuscripts. A more radical step to reducing re-review is to take revise-and-resubmit decisions off the table altogether [40]. Second, the scientific community can continue to explore ways of sharing reviews of rejected manuscripts, so that the associated review effort is not wasted. As our model with several journals suggests, much of the burden on the peer-review community flows from the need to solicit fresh reviews of already rejected manuscripts when those manuscripts are sent to new journals. Cascading review and open-review platforms represent promising moves in this direction.
@@ Acknowledgments
We thank Kyle Myers for bringing AO’s paper to our attention. This work was partially supported by NSF awards SES-2346645 to CTB and SES-2346644 to KG, and by TWCF Diverse Intelligences frameworks grant 32581 to CTB. KG thanks the University of Washington Department of Biology for visitor support and members of the Institute for Advanced Study in Toulouse for helpful discussions.
@@ A.2.1 Peer-review accuracy depends on reviewing load
Consider a single journal that sends every manuscript that it receives out for review. Let $L$ denote the review load, such that if the marginal author is $\tilde{q}$, then $L=1-\tilde{q}$. Let the review noise $\sigma_{Y}$ depend on $L$ through the function $\sigma_{Y}=\Sigma_{Y}(L)$. We assume $\Sigma_{Y}^{\prime}(L)\geq 0$. Write author $q$’s acceptance probability when facing review threshold $y$ with review noise $\sigma_{Y}$ as $a(q;y,\sigma_{Y})=\mathsf{Pr}\!\left\{Y\geq y|F_{X}(X)=q\right\}$ where $Y|q\sim\mathsf{N}(F_{X}^{-1}(q),\sigma^{2}_{Y})$. Writing equilibrium acceptance probabilities as $)\hat{a}(q)=a(q;\hat{y},\sigma_{Y}=\Sigma_{Y}(1-\hat{q}$)), the model equilibrium can again be found as the solution to AR and CF conditions:
[…]
$\int^{1}_{\hat{q}}\hat{a}(q)\,dq$ $=k.$
The dependence of $\sigma_{Y}$ on the review load is included via the new definition of $\hat{a}(q)$.
@@ A.2.2 Single journal with desk review
$\begin{pmatrix}D\\ Y\end{pmatrix}\sim\mathsf{N}_{2}\!\left(\begin{pmatrix}x/(1+\sigma^{2}_{X})\\ x/(1+\sigma^{2}_{X})\end{pmatrix},\dfrac{1}{1+\sigma^{2}_{X}}\begin{pmatrix}\sigma_{X}^{2}+\sigma_{D}^{2}+\sigma_{X}^{2}\sigma_{D}^{2}&\sigma^{2}_{X}\\ \sigma^{2}_{X}&\sigma_{X}^{2}+\sigma_{Y}^{2}+\sigma_{X}^{2}\sigma_{Y}^{2}\end{pmatrix}\right)$.
@@ A.3.3 Underuse of desk rejection from shared reliance on a common reviewer pool
If a journal sends a manuscript out for review, it must act in accord with the reviewer’s report. If the reviewer reports $Y=1$, the manuscript is published, otherwise the manuscript is rejected. The value to the journal of sending a manuscript to a reviewer with accuracy $(\alpha,\beta)$ is
$v(\alpha,\beta)$ $=v_{1}\mathsf{Pr}\!\left\{\theta=1,Y=1\right\}+v_{0}\mathsf{Pr}\!\left\{\theta=0,Y=1\right\}$
$=v_{1}(1-\beta)\rho+v_{0}\alpha(1-\rho).$
Note that the journal prefers to desk reject a manuscript instead of sending it to a reviewer whose report is independent of the manuscript’s suitability (i.e., $v(\alpha,1-\alpha)<0$ for any $\alpha$); the journal prefers to send a manuscript to a perfectly accurate reviewer instead of desk rejecting it ($v(0,0)>0$), and the journal is better off when the reviewer is more accurate (both $\partial v/\partial\alpha<0$ and $\partial v/\partial\beta<0$). Because a journal’s payoff will be based only on $v(\alpha,\beta)$ and not $\alpha$ and $\beta$ directly, label reviewers by the expected value that they generate for the journal, $v=v(\alpha,\beta)$, i.e., “reviewer vv”, and call $v$ the reviewer’s value.

### v3

@@ III.3.2 Several competing journals
A formal model that includes desk rejection with several competing journals lies outside the present scope, since it would require modeling how authors behave when they face journals with different desk-rejection cutoffs. However, intuition suggests that when competing journals rely on a common pool of reviewers, a “tragedy of the reviewer commons” [51] makes desk rejection less useful for managing the review load. The logic goes as follows. When a journal has its own private community of reviewers, it fully internalizes the cost of sending a manuscript our for review. At equilibrium, such a journal sets its desk-rejection policy to optimally trade off the cost of depleting its available review labor against the benefit of using peer review to learn more about a manuscript’s quality. But when several competing journals rely on the same common pool of reviewers, then a journal does not fully internalize the cost of requesting a review, because the cost of depleting the available review labor is borne by all journals. Yet the requesting journal still captures the full benefit of obtaining a review and learning more about the manuscript under consideration. Thus, because journals that rely on a shared reviewer pool do not fully internalize the cost of requesting reviews, they will set a desk-rejection policy that desk-rejects manuscripts less often than would be socially optimal.
@@ IV Discussion
Like all models, our model makes simplifications that provide scope for future work. Perhaps the most substantively, we treat manuscript quality as exogenous. Of course, authors are not endowed with fully formed manuscripts, but instead they decide how much effort to invest to developing manuscripts based on their anticipation of the journal’s scrutiny. For example, the rise of revise-and-resubmit as an author’s likely best outcome may perversely encourage authors to submit manuscripts that are less than polished, both because revision seems inevitable and because journals can no longer expect polished submissions if unpolished submissions become the norm [58]. A model of journal and author behavior that endogenizes manuscript quality would provide more satisfying insight on this count.
@@
If raising submission costs is off the table, journals might instead try to increase the supply of high-quality review labor. But journals that seek to recruit more high-quality reviewing are handcuffed by the tradition of peer review as a volunteer activity. In a typical labor market, employers can attract more labor by increasing wages. Journals operate without this basic market mechanism when reviewers are unpaid. If publishers were able to recruit more high-quality review labor by paying appropriate wages, they could maintain constant review quality in the face of an increase in submissions, breaking the feedback loop between screening and sorting.
@@
To be sure, paying reviewers brings risks. Non-profit journals would need to pass the cost of reviewer pay on to authors or subscribers, although perhaps society journals could maintain a base of volunteer reviewers by cultivating a sense of shared responsibility for the journal. Paying reviewers also commodifies review labor, likely irreversibly [59]. This commodification could perversely decrease the availability of high-quality review labor if reviewers operating under the honor system perceive their effort to be worth more than the wages that the journal offers [59, 26, 60]. Nevertheless, recent experiments with modest payments to reviewers have succeeded in accelerating the pace of peer review, at least in the short run [23, 25, 26]. Wages needn’t be simple direct payments, either. An intriguing alternative idea is to offer monetary prizes for top reviews, which could sweeten the pot for reviewers while motivating reviewer effort [61]. While substantial trial and error should be expected in establishing a plan that adequately compensates reviewers, once such a plan is established, it could give publishers a lever to recruit reviewers that they currently lack.
@@
Are there other options that might salvage the institution of (voluntary) peer review? We highlight a few options here, all of which echo the suggestions of others. First, journals can reduce re-review of revised manuscripts [62, 50]. Re-review is currently commonplace: accepted papers frequently go through at least two [63] or more [50] rounds of review prior to acceptance. But routinely sending revised manuscripts back out for review only burdens the reviewer community further while encouraging both authors and reviewers to modulate their effort in anticipation of the revision and re-review process to follow [58]. Journals can reduce re-review by empowering, or even encouraging, academic editors to be more proactive in their decision-making, especially when evaluating revised manuscripts. A more radical step to reducing re-review is to take revise-and-resubmit decisions off the table altogether [58]. Second, the scientific community can continue to explore ways of sharing reviews of rejected manuscripts, so that the associated review effort is not wasted. As our model with several journals suggests, much of the burden on the peer-review community flows from serial re-review of already rejected manuscripts when those manuscripts are sent to new journals. Cascading review and open-review platforms represent promising moves in this direction.
Finally, one last lever that our model suggests for salvaging peer review is to decrease $v$, the reward that authors reap for publishing in selective journals. In this model, we have taken the view that $v$ is set exogeneously by the scientific community and, unlike the other solutions we have considered, cannot be changed unilaterally by any single actor. Further, the forces that determine $v$ are varied and complex because the rewards attached to publication serve a variety of economic and sociological functions in science. For example, we [64] and others [65] have argued that rewarding publication constitutes a rational if imperfect scheme for motivating researchers to work hard and to take scientific risks while also protecting researchers’ livelihoods from the vicissitudes of scientific chance. But this work does not consider the pressure that publication rewards place on peer review. Perhaps a more holistic view would suggest that the scientific endeavor would be better off if fewer rewards flowed to authors who publish in top outlets. Or perhaps it wouldn’t; the answer remains far from clear. Regardless, whether science would be better off if the rewards to publication were reduced, and if so, how the scientific community might coordinate to change the prestige associated with publication seem to be rich and increasingly urgent topics to contemplate.
@@ Acknowledgments
Having written a paper about the demise of peer review, we were heartened to receive three thoughtful, detailed, and on-point reviews of this manuscript. We thank the reviewers for their careful attention and constructive feedback, and for demonstrating that all is not yet lost. We also thank Kyle Myers and Nihar Shah for bringing AO’s and Zhang et al.’s papers to our attention. This work was partially supported by NSF awards SES-2346645 to CTB and SES-2346644 to KG, and by TWCF Diverse Intelligences frameworks grant 32581 to CTB. KG thanks both the Department of Biology at the University of Washington and the Institute for Advanced Study in Toulouse for visitor support, the latter via funding from the French National Research Agency (ANR) under grant ANR-17-EURE-0010 (Investissements d’Avenir program).
@@ A.1 Proofs
We make two notes about the proof above. First, as AO emphasize, the effect of increasing review noise on the journal’s cutoff $\hat{y}$ is ambiguous. However, Lehmann’s condition guarantees that, regardless of whether $\hat{y}$ increases or decreases, the acceptance probability $\hat{a}_{1}(q)<\hat{a}(q)$ for all $q>\hat{q}$, which drives the main result. Second, the applicability of Lehmann’s condition depends on the details of the conditional distribution of $Y$ given $X$ (or alternatively the conditional distribution of $Y$ given $q$). In our model, Lehmann’s condition follows immediately from the fact that $(X,Y)$ has a bivariate normal distribution (which itself follows from our assumption that the triple $(\theta,X,Y)$ has a trivariate normal distribution). As AO make clear, Lehmann’s condition, and hence the above proof, holds more broadly; indeed, AO show that for Lehmann’s condition to hold it suffices that the conditional distribution of $Y$ given $X$ has a location-scale structure $Y=X+\sigma_{Y}\epsilon$ for some random variable $\epsilon$. However, Lehmann’s condition does not necessarily hold universally for all distributions of $(\theta,X,Y)$. Articulating exactly the collection of distributions on $(\theta,X,Y)$ for which Lehmann’s condition holds is beyond the scope of this article.
@@ A.2.1 Peer-review accuracy depends on reviewing load
Consider a single journal that sends every manuscript that it receives out for review. Let $L$ denote the review load, such that if the marginal author is $\tilde{q}$, then $L=1-\tilde{q}$. Let the review noise $\sigma_{Y}$ depend on $L$ through the function $\sigma_{Y}=\Sigma_{Y}(L)$. We assume $\Sigma_{Y}^{\prime}(L)\geq 0$. Write author $q$’s acceptance probability when facing review threshold $y$ with review noise $\sigma_{Y}$ as $a(q;y,\sigma_{Y})=\mathsf{Pr}\!\left\{Y\geq y|F_{X}(X)=q\right\}$ where $Y|q\sim\mathsf{N}(F_{X}^{-1}(q),\sigma^{2}_{Y})$. Writing equilibrium acceptance probabilities as $)\hat{a}(q)=a(q;\hat{y},\sigma_{Y}=\Sigma_{Y}(1-\hat{q}$)), the model equilibrium can again be found as the solution to AR and CF conditions: {align*} v ^a(^q) = c
∫^1_^q ^a(q) dq = k. The dependence of $\sigma_{Y}$ on the review load is included via the new definition of $\hat{a}(q)$.
@@ A.2.2 Single journal with desk review
[…]
@@ A.3.3 Underuse of desk rejection from shared reliance on a common reviewer pool
If a journal sends a manuscript out for review, it must act in accord with the reviewer’s report. If the reviewer reports $Y=1$, the manuscript is published, otherwise the manuscript is rejected. The value to the journal of sending a manuscript to a reviewer with accuracy $(\alpha,\beta)$ is {align*} v(α, β) = v_1 $\mathsf{Pr}\!\left\{\theta=1,Y=1\right\}$ + v_0 $\mathsf{Pr}\!\left\{\theta=0,Y=1\right\}$
= v_1 (1 - β) ρ+ v_0 α(1 - ρ). Note that the journal prefers to desk reject a manuscript instead of sending it to a reviewer whose report is independent of the manuscript’s suitability (i.e., $v(\alpha,1-\alpha)<0$ for any $\alpha$); the journal prefers to send a manuscript to a perfectly accurate reviewer instead of desk rejecting it ($v(0,0)>0$), and the journal is better off when the reviewer is more accurate (both $\partial v/\partial\alpha<0$ and $\partial v/\partial\beta<0$). Because a journal’s payoff will be based only on $v(\alpha,\beta)$ and not $\alpha$ and $\beta$ directly, label reviewers by the expected value that they generate for the journal, $v=v(\alpha,\beta)$, i.e., “reviewer vv”, and call $v$ the reviewer’s value.